OpenStudy (anonymous):

a 1.00 kg block compresses a spring for which K1 = 100 N/m by 20.0 cm and is then released to move across a horizontal, frictionless table, where it hits and compresses another spring for which K2 = 50 N/m. Determine: a) the total mechanical energy of the system b) the speed of the mass while moving freely between springs, and c) the maximum compression of the second spring

OpenStudy (alfie):

 Alright this is the situation. In the initial situation we have the following... \[P_e = \frac{1}{2}kx^2\] And that's pretty much it, let's we see in the moment when the spring lets go the block. All the Potential energy of the spring is transformed to kinetic energy. \[\frac{1}{2}kx^2 = \frac{1}{2}mv^2\] This way you can solve for v and get the velocity.